# Confronting Cold New Early Dark Energy and its Equation of State with Updated CMB and Supernovae Data

### v1

@@ Abstract
Cold New Early Dark Energy (Cold NEDE) addresses the Hubble tension through a triggered vacuum phase transition in the dark sector. In this paper, we constrain a phenomenological fluid model using recent cosmic microwave background likelihoods based on Planck NPIPE data alongside baryonic acoustic oscillations and supernovae data from Pantheon+. Exploiting the enhanced constraining power of the datasets, we introduce and study an extended version of the NEDE fluid model in which the equation of state parameter $w_{\mathrm{NEDE}}$, characterizing the post-phase transition fluid, is allowed to evolve with non-vanishing derivatives ${d}w_{\mathrm{NEDE}}/d\ln a$ and ${d^{2}}w_{\mathrm{NEDE}}/{d}(\ln a)^{2}$. Our results indicate that data is compatible with a rather simple time dependence that could arise from a mixture of radiation and a stiff fluid. With the updated datasets, the base and extended models still show a significant reduction of the DMAP tension from $6.3\sigma$ in $\Lambda$CDM down to $3.5\sigma$ with a small simultaneous reduction of the $S_{8}$ tension, slightly improving over recent findings for the axion-like early dark energy model.
@@ I Introduction
Since Cold NEDE, like other EDE-type models, fit the CMB with a much lower sound horizon and larger $H_{0}$, it is really a significantly different fit to the CMB than $\Lambda$CDM. In (N)EDE models the lower sound horizon and the increased expansion rate due to the presence of the (N)EDE component is accompanied by a shortening of the CMB damping scale and a faster decay of the Weyl potential due to new acoustic oscillations in the (N)EDE fluid, which have to be compensated by a bluer primordial spectral index $n_{s}$ and a higher amplitude of primordial perturbation $A_{s}$, as well as a larger fraction of Cold Dark Matter (CDM) $\omega_{cdm}$, respectively. The different predictions of the cosmological parameters and of the details of the CMB TT, TE and EE power spectrum as well as of the matter power spectrum depending on the details of the specific (N)EDE model, leads to the expectation that more precise CMB and BAO data will on their own be able to single out the right model beyond $\Lambda$CDM. Here, we will take one more step in this direction by testing Cold NEDE against $\Lambda$CDM with the improved CMB data based on the NPIPE maps [25], and compare with an earlier similar test of AxiEDE also using NPIPE data [26]. On top of a better handle of a number of systematic effects, the new NPIPE maps have a larger signal-to-noise at small scales thanks to improvements in the processing of time ordered data [25], allowing to use a larger sky fraction (about $\sim 80\%$) and resulting in $\sim 10\%$ higher precision on $\Lambda$CDM parameters [27, 28]. In addition, the Pantheon dataset used in previous studies of the Cold NEDE model is updated with Pantheon+ [29].
@@
The paper is organized as follows. In Sec. II, we review the Cold NEDE model and introduce the extended model with the improved NEDE fluid treatment. Then in Sec. III, we discuss the data analysis and the results, and then finally in Sec. IV, we conclude.
@@ II.1 Minimal Phenomenological Setup
$\delta\phi^{\prime\prime}+2H\delta\phi^{\prime}+(k^{2}/a^{2}+m^{2})\delta\phi$ $=-\frac{1}{2}h^{\prime}\phi^{\prime}\,,$ (4b)
where we switched to momentum space, a prime denotes differentiation with respect to conformal time $\tau$, and $h$ denotes the spatial trace of the metric perturbation. For $H\gg m$, the background equation describes an overdamped harmonic oscillator with $\bar{\phi}\simeq\bar{\phi}_{\mathrm{ini}}$. As $H$ drops below the trigger mass $m$, the field thaws and starts to roll down to the bottom of its potential where it oscillates with frequency $m$. The transition in $w_{\mathrm{NEDE}}$ is triggered during the thawing phase when a threshold value $\phi_{*}$ is reached. This happens before the field reaches the minimum of its potential for the first time, corresponding to $H(z_{*})\sim m$. The precise value is determined by the microscopic model. For example, in the case of Cold NEDE it is [18] $H(z_{*})/m\simeq 0.2$. This condition implicitly fixes the transition redshift $z_{*}$ (and hence also $\phi_{*}$). After it has served its main purpose by triggering the transition in $w_{\mathrm{NEDE}}$, due to its coherent oscillations, $\phi$ acts like a fuzzy dark matter component [30, 31, 32], where its abundance today is controlled by $\phi_{\mathrm{ini}}$ [19],
@@
$c_{s}^{2}=w_{\mathrm{NEDE}}(t)-\frac{1}{3}\frac{w_{\mathrm{NEDE}}^{\prime}(\tau)}{1+w_{\mathrm{NEDE}}(\tau)}\frac{1}{\mathcal{H}}\,.$ (9)
This is a crucial difference compared to AxiEDE, which can be described in terms of fluid perturbations with a $k$-dependent sound speed [38] (wheras $c_{s}$ here is $k$-independent). Moreover, in the case of the phenomenological NEDE model, the fluid perturbations are initialized through the $k$-dependent matching equations in (7). In particular, the choice of the trigger plays an important role in determining the subsequent dynamics. This was recently demonstrated in the case of Hot NEDE, where the ultralight field was replaced by a dark sector temperature [22]. In any event, this trigger sector in the NEDE model does not have a counterpart in AxiEDE.
@@ III.1 Implementation and Datasets
The phenomenological NEDE model, implemented in the Boltzmann code TriggerCLASS [^3] We use TriggerCLASS version 6.1, which is publicly available: https://github.com/NEDE-Cosmo/TriggerCLASS.git and modifies CLASS [45]., is interfaced with the MCMC sampler MontePython [46, 47]. We scan the six $\Lambda$CDM parameters ($\omega_{b}$, $\omega_{\mathrm{cdm}}$, $h$, $\log_{10}(A_{s})$, $n_{s}$, $\tau_{\rm reio}$) using the Metropolis-Hastings algorithm with standard prior ranges. The neutrino sector is modeled with two massless and one massive neutrino with mass $m_{3}=0.06\mathrm{eV}$ and temperature $T_{3}=0.716\mathrm{eV}$, amounting to $N_{\mathrm{eff}}=3.046$. In accordance with the discussion in Sec. II, we sample the four NEDE parameters $f_{\mathrm{NEDE}}\in[0,0.3]$, $\log_{10}(z_{*})\in[2.8,4]$, $3w\in[1,3]$, and $\Omega_{\phi}\in[0,0.1]$ when testing the NEDE model in Sec. III.2. As explained in more detail in Sec. III.3, in the case of the extended model, we add additional parameters allowing for a non-trivial time evolution of the equation of state parameter after the transition [see also (10)].
@@
  • BAO: Baryonic acoustic oscillation (BAO) measurements [^4] It is worth mentioning that new data from DESI has been published as this work was under completion [50]. Given that DESI data shows some hints of deviations from $\Lambda$CDM that goes in the direction of relaxing constraints on EDE models [51, 52], our expectation is that making use of SDSS data (that have roughly the same statistical power) leads to somewhat stronger constraints. We defer to future work to test that explicitly. from the 6dF Galaxy Survey [53] (at $z=0.106$), SDSS DR7 [54] (at $z=0.15$), and CMASS and LOWZ galaxy samples of BOSS DR12 [55] (at $z=0.38$, $0.51$, $0.61$). The latter includes a constraint on the growth rate $f\sigma_{8}$ from redshift space distortions.
@@
Figure 1: Comparison of 2D posterior distributions reconstructed from either the CamSpec/NPIPE, HiLLiPoP/NPIPE, or Plik/Planck 2018 likelihoods (alongside the baseline datasets) in the base NEDE model. Here and henceforth, the horizontal and vertical grey bands show the $1\sigma$ and $2\sigma$ contours of the constraints on $S_{8}$ from [56] and on $H_{0}$ from [1]. Compared to Planck 2018, both NPIPE likelihoods lead to tighter constraints on $f_{\mathrm{NEDE}}$ and a later decay (closer to matter-radiation equality). The model’s ability to address the $S_{8}$ tension is preserved for CamSpec/NPIPE and even enhanced for HiLLiPoP/NPIPE.
@@
It is a quantitative estimate of the inconsistency between the SH0ES measurement and our baseline datasets consisting of CMB, BAO, and Pantheon+. The benefit of this method is that it is insensitive to prior volume effects, which are known to make the Gaussian tension less reliable for EDE-type models [58, 24]. We use the simulated-annealing optimizer Procoli [59] to find the global maximum likelihoods value reported in our data tables. All chains satisfy a Gelman-Rubin convergence level of $|R-1|\lesssim 0.05$ [60]. When plotting marginalized posteriors, we show the $68\%$ C.L. and $95\%$ C.L contours.
@@ III.2 Confronting NEDE with NPIPE (and Pantheon+)
Plik/Planck 2018
SH0ES prior? no yes
$h$ 0.6867($0.6867(0.6913)^{+0.0071}_{-0.013}$ $0.7205(0.7216)\pm 0.0079$
$f_{\mathrm{NEDE}}$ $<0.119(0.068)$ 0.157($0.157(0.162)^{+0.026}_{-0.023}$
$\log_{10}(z_{*})$ 3.67($3.67(3.72)^{+0.22}_{-0.13}$ 3.662($3.662(3.678)^{+0.065}_{-0.055}$
$\Omega_{\phi}$ $<0.00629(0.0003)$ 0.0038($0.0038(0.0032)^{+0.0014}_{-0.0033}$
$3w_{*}$ $2.10(1.98)^{+0.37}_{-0.48}$ 2.058($2.058(2.04)^{+0.098}_{-0.14}$
$10^{2}\omega{}_{b}$ 2.261($2.261(2.266)^{+0.019}_{-0.026}$ $2.307(2.311)\pm 0.024$
$\omega{}_{\rm cdm}$ 0.1234($0.1234(0.1257)^{+0.0022}_{-0.0040}$ $0.1325(0.1331)\pm 0.0030$
$n_{s}$ 0.9718($0.9718(0.9769)^{+0.0057}_{-0.0080}$ $0.9909(0.9934)\pm 0.0056$
$10^{9}A_{s}$ 2.126($2.126(2.123)^{+0.029}_{-0.035}$ 2.163($2.163(2.157)^{+0.029}_{-0.033}$
$\tau{}_{\rm reio}$ 0.0569($0.0569(0.0548)^{+0.0067}_{-0.0076}$ 0.0580($0.0580(0.0565)^{+0.0068}_{-0.0077}$
$\Omega_{m}$ $0.3122(0.3107)\pm 0.0058$ $0.3034(0.3032)\pm 0.0053$
$S_{8}$ 0.822($0.822(0.84)^{+0.017}_{-0.014}$ 0.823($0.823(0.828)^{+0.022}_{-0.019}$
$\Delta\chi^{2}({\rm NEDE}-\Lambda$CDM) $-1.94$ $-33.34$
@@
Table 2: Confidence intervals and $1\sigma$ error bars (best-fit in parenthesis) for the base NEDE model reconstructed from the Plik/Planck 2018 likelihood together with our baseline datasets, with and without the SH0ES prior on $M_{b}$.
CamSpec/NPIPE HiLLiPoP/NPIPE Plik/Planck 2018
Planck $-\ell$ TTTEEE 11237.56 11245.36 30494.32 30501.97 2344.87 2350.34
Planck $-\ell$ TT 22.61 21.39 21.95 21.04 21.98 20.73
Planck $-\ell$ EE 396.07 396.41 -- -- 396.09 396.28
Lollipop EB -- -- 125.66 126.78 -- --
Planck lensing 9.27 9.84 9.13 9.42 9.1 9.85
@@
BOSS BAO/$f\sigma_{8}$ DR12 6.82 6.04 5.91 6.00 7.19 6.22
Pantheon+ 1411.06 1413.16 1411.08 1412.87 1411.18 1412.4
SH0ES -- 2.66 -- 3.91 -- 2.27
total $\chi^{2}_{\rm min}$ 13084.52 13096.83 32070.27 32083.69 4191.52 4199.21
$\Delta\chi^{2}_{\rm min}({\rm NEDE}-\Lambda{\rm CDM})$ $-1.44$ $-28.54$ $-2.4$ $-30.04$ $-1.94$ $-33.34$
@@
Table 3: Best-fit $\chi^{2}$ per experiment (and total) for the base NEDE model when fitted to our baseline datasets supplemented either by the CamSpec/NPIPE, HiLLiPoP/NPIPE or Plik/Planck 2018 likelihood. We compare the fits with and without the SH0ES prior on $M_{b}$. We report the $\Delta\chi^{2}_{\rm min}\equiv\chi^{2}_{\rm min}({\rm NEDE})-\chi^{2}_{\rm min}(\Lambda{\rm CDM})$ and the tension metric $Q_{\rm DMAP}$ (for reference, $Q_{\mathrm{DMAP}}$ values for $\Lambda$CDM are also shown.)

### v2

# Confronting Cold New Early Dark Energy and its Equation of State with Updated CMB, Supernovae, and BAO Data

@@ Abstract
Cold New Early Dark Energy (Cold NEDE) addresses the Hubble tension through a triggered vacuum phase transition in the dark sector. In this paper, we constrain a phenomenological fluid model using recent cosmic microwave background likelihoods based on Planck NPIPE data alongside baryonic acoustic oscillations (BAO) and supernovae data from Pantheon+. Exploiting the enhanced constraining power of the datasets, we introduce and study an extended version of the NEDE fluid model in which the equation of state parameter $w_{\mathrm{NEDE}}$, characterizing the post-phase transition fluid, is allowed to evolve with non-vanishing derivatives ${d}w_{\mathrm{NEDE}}/d\ln a$ and ${d^{2}}w_{\mathrm{NEDE}}/{d}(\ln a)^{2}$. Our results indicate that data is compatible with a rather simple time dependence that could arise from a mixture of radiation and a stiff fluid. With the updated datasets, the base and extended models still show a significant reduction of the DMAP tension from $6.3\sigma$ in $\Lambda$CDM down to $3.5\sigma$ with a small simultaneous reduction of the $S_{8}$ tension, slightly improving over recent findings for the axion-like early dark energy model. Finally, we also provide a first test of the model against new BAO data from the Dark Energy Spectroscopic Instrument (DESI) survey. Replacing the previous BAO constraints in our analysis with the new ones, the tension is further reduced to $2.6\sigma$, reaffirming the Cold NEDE model as a promising solution to the Hubble tension.
@@ I Introduction
Since Cold NEDE, like other EDE-type models, fit the CMB with a much lower sound horizon and larger $H_{0}$, it is really a significantly different fit to the CMB than $\Lambda$CDM. In (N)EDE models the lower sound horizon and the increased expansion rate due to the presence of the (N)EDE component is accompanied by a shortening of the CMB damping scale and a faster decay of the Weyl potential due to new acoustic oscillations in the (N)EDE fluid, which have to be compensated by a bluer primordial spectral index $n_{s}$ and a higher amplitude of primordial perturbation $A_{s}$, as well as a larger fraction of Cold Dark Matter (CDM) $\omega_{cdm}$, respectively. The different predictions of the cosmological parameters and of the details of the CMB TT, TE and EE power spectrum as well as of the matter power spectrum depending on the details of the specific (N)EDE model, leads to the expectation that more precise CMB and BAO data will on their own be able to single out the right model beyond $\Lambda$CDM. Here, we will take one more step in this direction by testing Cold NEDE against $\Lambda$CDM with the improved CMB data based on the NPIPE maps [25], and compare with an earlier similar test of AxiEDE also using NPIPE data [26]. On top of a better handle of a number of systematic effects, the new NPIPE maps have a larger signal-to-noise at small scales thanks to improvements in the processing of time ordered data [25], allowing to use a larger sky fraction (about $\sim 80\%$) and resulting in $\sim 10\%$ higher precision on $\Lambda$CDM parameters [27, 28]. In addition, the Pantheon dataset used in previous studies of the Cold NEDE model is updated with Pantheon+ [29]. We also provide a first test of the model against new BAO data from the Spectroscopic Dark Energy Instrument (DESI) [30]. To that end, we perform an analysis where we replace our baseline BAO datasets with the new constraints. Interestingly, DESI data has previously shown some hints of deviations from $\Lambda$CDM that go in the direction of relaxing constraints on EDE models [31, 32].
@@
The paper is organized as follows. In Sec. II, we review the Cold NEDE model and introduce the extended model with the improved NEDE fluid treatment. Then in Sec. III, we discuss the data analysis and the results, and then in Sec. IV, we conclude. In Appendix A, we provide likelihood profiles to further support our statistical findings.
@@ II.1 Minimal Phenomenological Setup
$\delta\phi^{\prime\prime}+2H\delta\phi^{\prime}+(k^{2}/a^{2}+m^{2})\delta\phi$ $=-\frac{1}{2}h^{\prime}\bar{\phi}^{\prime}\,,$ (4b)
where we switched to momentum space, a prime denotes differentiation with respect to conformal time $\tau$, and $h$ denotes the spatial trace of the metric perturbation. For $H\gg m$, the background equation describes an overdamped harmonic oscillator with $\bar{\phi}\simeq\bar{\phi}_{\mathrm{ini}}$. As $H$ drops below the trigger mass $m$, the field thaws and starts to roll down to the bottom of its potential where it oscillates with frequency $m$. The transition in $w_{\mathrm{NEDE}}$ is triggered during the thawing phase when a threshold value $\phi_{*}$ is reached. This happens before the field reaches the minimum of its potential for the first time, corresponding to $H(z_{*})\sim m$. The precise value is determined by the microscopic model. For example, in the case of Cold NEDE it is [18] $H(z_{*})/m\simeq 0.2$. This condition implicitly fixes the transition redshift $z_{*}$ (and hence also $\phi_{*}$). After it has served its main purpose by triggering the transition in $w_{\mathrm{NEDE}}$, due to its coherent oscillations, $\phi$ acts like a fuzzy dark matter component [33, 34, 35] (see also [36, 37]), where its abundance today is controlled by $\phi_{\mathrm{ini}}$ [19],
@@
$c_{s}^{2}=w_{\mathrm{NEDE}}(\tau)-\frac{1}{3}\frac{w_{\mathrm{NEDE}}^{\prime}(\tau)}{1+w_{\mathrm{NEDE}}(\tau)}\frac{1}{\mathcal{H}}\,.$ (9)
This is a crucial difference compared to AxiEDE, which deep inside its dissipating phase can be described in terms of a fluid with a $k$-dependent sound speed [43] (whereas $c_{s}$ here is $k$-independent). Moreover, in the case of the phenomenological NEDE model, the fluid perturbations are initialized through the $k$-dependent matching equations in (7). In particular, the choice of the trigger plays an important role in determining the subsequent dynamics. This was recently demonstrated in the case of Hot NEDE, where the ultralight field was replaced by a dark sector temperature [22], which led to a decisively different phenomenology. For Cold NEDE, the parameter combination $H(z_{*})/m$, which is predicted by the microscopic model, controls how much the trigger field has thawed by the time of the phase transition. As a result, it determines $\bar{\phi}^{\prime}$ on the right hand side of the matching equations 7 and thus controls the initial size of the perturbations $\delta_{\mathrm{NEDE}}(\tau_{*},\mathbf{k})$ in the dissipating dark sector fluid. As we will see in Sec. III.4, this leads to observable consequences. In any event, this trigger sector in the NEDE model does not have a counterpart in AxiEDE.
@@ III.1 Implementation and Datasets
The phenomenological NEDE model, implemented in the Boltzmann code TriggerCLASS [^3] We use TriggerCLASS version 6.1, which is publicly available: https://github.com/NEDE-Cosmo/TriggerCLASS.git and modifies CLASS [50]., is interfaced with the MCMC sampler MontePython [51, 52]. We scan the six $\Lambda$CDM parameters ($\omega_{b}$, $\omega_{\mathrm{cdm}}$, $h$, $\log_{10}(A_{s})$, $n_{s}$, $\tau_{\rm reio}$) using the Metropolis-Hastings algorithm with standard prior ranges. The neutrino sector is modeled with two massless and one massive neutrino with mass $m_{3}=0.06\mathrm{eV}$ and temperature $T_{3}=0.716\mathrm{eV}$, amounting to $N_{\mathrm{eff}}=3.046$. In accordance with the discussion in Sec. II, we sample the four NEDE parameters [^4] We checked for our base model runs with NPIPE data that relaxing the upper prior boundary did not lead to any relevant changes in the parameter inference. $f_{\mathrm{NEDE}}\in[0,0.3]$, $\log_{10}(z_{*})\in[2.8,4]$, $3w\in[1,3]$, and $\Omega_{\phi}\in[0,0.1]$ when testing the NEDE model in Sec. III.2. As explained in more detail in Sec. III.3, in the case of the extended model, we add additional parameters allowing for a non-trivial time evolution of the equation of state parameter after the transition [see also (10)].
@@
  • BAO: Baryonic acoustic oscillation (BAO) measurements from the 6dF Galaxy Survey [55] (at $z=0.106$), SDSS DR7 [56] (at $z=0.15$), and CMASS and LOWZ galaxy samples of BOSS DR12 [57] (at $z=0.38$, $0.51$, $0.61$). The latter includes a constraint on the growth rate $f\sigma_{8}$ from redshift space distortions.
@@
Figure 1: Comparison of 2D posterior distributions reconstructed from either the CamSpec/NPIPE, HiLLiPoP/NPIPE, or Plik/Planck 2018 likelihoods (alongside our baseline datasets) in the base NEDE model. The green contours are obtained from an analysis where our baseline BAO data has been exchanged with DESI data. Here and henceforth, the horizontal and vertical gray bands show the $1\sigma$ and $2\sigma$ constraints on $S_{8}$ from [58] and on $H_{0}$ from [1]. Compared to Planck 2018, both NPIPE likelihoods lead to tighter constraints on $f_{\mathrm{NEDE}}$ and a later decay (closer to matter-radiation equality). The model’s ability to address the $S_{8}$ tension is preserved for CamSpec/NPIPE and even enhanced for HiLLiPoP/NPIPE.
We will refer to BAO and Pantheon+ as our baseline dataset combination. In this work, we also explore the impact of the BAO measurements from the recently released Dark Energy Spectroscopic Instrument (DESI) survey [30]. More specifically, for the runs labeled “DESI” we replace the BAO baseline dataset from above with DESI BAO data. To that end, we employ the same likelihood that has recently been used for constraining the axiEDE model in [32] and was build from the data in Tab. 1 of [30].
@@
Here, $\chi^{2}$ represents the best-fit chi-squared with ($\mathrm{w\backslash\,H_{0}}$) and without ($\mathrm{w\backslash o\,H_{0}}$) including the SH0ES measurement in the analysis. The DMAP tension is a quantitative estimate of the inconsistency between the SH0ES measurement and the other datasets consisting of CMB, BAO, and Pantheon+ as described above. For Gaussian posteriors this measure is equivalent to the usual tension measure computed from the mean values and their confidence intervals. In particular if the two dataset combinations are consistent, their maximum likelihoods should be close, resulting in a small $Q_{\mathrm{DMAP}}$. In [60] the DMAP measure was used to compare different models in their attempt to resolve the Hubble tension. The benefit of this method is that it is insensitive to prior volume effects, which are known to make the Gaussian tension less reliable for EDE-type models [61, 24]. We use the simulated-annealing optimizer Procoli [62] to find the global maximum likelihoods value reported in our data tables. All chains satisfy a Gelman-Rubin convergence level of $|R-1|\lesssim 0.05$ [63]. When plotting marginalized posteriors, we show the $68\%$ C.L. and $95\%$ C.L. contours. Finally, in order to assess the overall fit quality of a given model, we employ the Akaike Information Criterium (see also [60]). It assigns a penalty of $2$ to the usual $\chi^{2}$ measure for each free parameter the model adds to the six $\Lambda$CDM parameters. More concretely, $\Delta{\rm AIC}=\chi^{2}_{\rm min}({\rm NEDE})-\chi^{2}(\Lambda{\rm CDM})+2(N_{p,\mathrm{NEDE}}-N_{p,\Lambda{\rm CDM}}),$ where $N_{p,\Lambda{\rm CDM}}=6$ and $N_{p,\mathrm{NEDE}}$ is the number of free parameters in a given NEDE model ($N_{p,\mathrm{NEDE}}=10$ for the base model).
@@ III.2 Confronting NEDE with NPIPE (and Pantheon+)
Plik/Planck 2018 CamSpec/NPIPE (DESI)
SH0ES prior? no yes no yes
$h$ 0.6867($0.6867(0.6913)^{+0.0071}_{-0.013}$ $0.7205(0.7216)\pm 0.0079$ 0.6886($0.6886(0.6839)^{+0.0070}_{-0.012}$ $0.7223(0.7225)\pm 0.0080$
$f_{\rm NEDE}$ $<0.119(0.068)$ 0.157($0.157(0.162)^{+0.026}_{-0.023}$ $<0.109(0.019)$ $0.146(0.145)\pm 0.024$
$\log_{10}(z_{*})$ 3.67($3.67(3.72)^{+0.22}_{-0.13}$ 3.662($3.662(3.678)^{+0.065}_{-0.055}$ unconstrained $(4.0)$ 3.570($3.570(3.585)^{+0.051}_{-0.065}$
$\Omega_{\phi}$ $<0.00629(0.0003)$ 0.0038($0.0038(0.0032)^{+0.0014}_{-0.0033}$ $<0.00497(0.00008)$ $<0.00589(0.00011)$
$3w_{*}$ $2.10(1.98)^{+0.38}_{-0.48}$ 2.058($2.058(2.04)^{+0.098}_{-0.14}$ 1.99($1.99(1.22)^{+0.44}_{-0.58}$ 2.11($2.11(2.09)^{+0.13}_{-0.20}$
$10^{-2}\omega{}_{b}$ 2.261($2.261(2.266)^{+0.019}_{-0.026}$ $2.307(2.311)\pm 0.024$ 2.234($2.234(2.229)^{+0.016}_{-0.020}$ 2.261($2.261(2.254)^{+0.019}_{-0.021}$
$\omega{}_{\rm cdm}$ 0.1234($0.1234(0.1257)^{+0.0022}_{-0.0040}$ $0.1325(0.1331)\pm 0.0030$ 0.1217($0.1217(0.1209)^{+0.0019}_{-0.0035}$ $0.1304(0.1309)\pm 0.0027$
$n_{s}$ 0.9718($0.9718(0.9769)^{+0.0057}_{-0.0080}$ $0.9909(0.9934)\pm 0.0056$ 0.9718($0.9718(0.9738)^{+0.0056}_{-0.0071}$ $0.9878(0.9899)\pm 0.0055$
$10^{9}A_{s}$ 2.126($2.126(2.123)^{+0.029}_{-0.035}$ 2.163($2.163(2.157)^{+0.029}_{-0.033}$ 2.124($2.124(2.113)^{+0.031}_{-0.035}$ 2.160($2.160(2.151)^{+0.030}_{-0.034}$
$\tau{}_{\rm reio}$ 0.0569($0.0569(0.0548)^{+0.0067}_{-0.0076}$ 0.0580($0.0580(0.0565)^{+0.0068}_{-0.0077}$ $0.0592(0.0577)\pm 0.0075$ 0.0596($0.0596(0.0578)^{+0.0066}_{-0.0079}$
$\Omega_{m}$ $0.3122(0.3107)\pm 0.0058$ $0.3034(0.3032)\pm 0.0053$ $0.3056(0.3063)\pm 0.0053$ $0.2956(0.2941)\pm 0.0047$
$S_{8}$ 0.822($0.822(0.84)^{+0.017}_{-0.014}$ 0.823($0.823(0.828)^{+0.022}_{-0.019}$ 0.815($0.815(0.825)^{+0.015}_{-0.012}$ 0.816($0.816(0.834)^{+0.019}_{-0.014}$
$\Delta\chi^{2}({\rm NEDE}-\Lambda$CDM) $-1.94$ $-33.34$ $-2.5$ $-33.42$
@@
Table 2: Confidence intervals and $1\sigma$ error bars (best-fit in parenthesis) for the base NEDE model reconstructed from the Plik/Planck 2018 and CamSpec/NPIPE likelihood together with our baseline datasets. Both analyses where performed with and without the SH0ES prior on $M_{b}$. In the case of NPIPE, the BAO data is taken from the DESI survey resulting in the smalles residual tension.
CamSpec/NPIPE HiLLiPoP/NPIPE Plik/Planck 2018 CamSpec/NPIPE (DESI)
SH0ES prior? no yes no yes no yes no yes
Planck $-\ell$ TTTEEE 11237.56 11245.36 30494.32 30501.97 2344.87 2350.34 11238.2 11244.44
Planck $-\ell$ TT 22.61 21.39 21.95 21.04 21.98 20.73 22.02 20.98
Planck $-\ell$ EE 396.07 396.41 -- -- 396.09 396.28 396.73 396.61
Lollipop EB -- -- 125.66 126.78 -- -- -- --
Planck lensing 9.27 9.84 9.13 9.42 9.1 9.85 9.54 9.83
@@
BOSS BAO/$f\sigma_{8}$ DR12 6.82 6.04 5.91 6.00 7.19 6.22 -- --
DESI BAO -- -- -- -- -- -- 16.13 12.92
Pantheon+ 1411.06 1413.16 1411.08 1412.87 1411.18 1412.4 1411.81 1414.36
SH0ES -- 2.66 -- 3.91 -- 2.27 -- 2.34
total $\chi^{2}_{\rm min}$ 13084.52 13096.83 32070.27 32083.69 4191.52 4199.21 13094.44 13101.48
$\Delta\chi^{2}_{\rm min}({\rm NEDE}-\Lambda{\rm CDM})$ $-1.44$ $-28.54$ $-2.4$ $-30.04$ $-1.94$ $-33.34$2.5 -33.42
@@
AIC 6.56 -20.54 5.6 -22.04 6.06 -25.34 5.5 -25.42
Table 3: Best-fit $\chi^{2}$ per experiment (and total) for the base NEDE model when fitted to our baseline datasets supplemented either by the CamSpec/NPIPE, HiLLiPoP/NPIPE or Plik/Planck 2018 likelihood. In the fourth column with CamSpec/NPIPE, the BAO data is taken from the DESI survey. We compare the fits with and without the SH0ES prior on $M_{b}$. We report the $\Delta\chi^{2}_{\rm min}\equiv\chi^{2}_{\rm min}({\rm NEDE})-\chi^{2}_{\rm min}(\Lambda{\rm CDM})$, the tension metric $Q_{\rm DMAP}$ and the value of $\Delta{\rm AIC}$ (for reference, $Q_{\mathrm{DMAP}}$ values for $\Lambda$CDM are also shown.)